OpenStudy (anonymous):

HALF LIFE PROBLEM The amount of a drug in a person's bloodstream can be modeled with exponential decay. Suppose that, in 3 hours, 18% of the drug is removed from the bloodstream. What is the half-life of the drug? A) 8.1 hours B) 8.7 hours C) 9.3 hours D) 9.9 hours E) 10.5 hours

OpenStudy (anonymous):

i learnt abou this not to long ago but i cant remember wat my answer was but i will have a think now an if i no ill write back :)

OpenStudy (anonymous):

Think of the half-life (in hours) as the reciprocal (because kind of the opposite, yes?) of the number of halvings in an hour. How many of these? How about: a third of the number of halvings in 3 hours. How many halvings in 3 hours? How about: the power you need to raise a half to, so it's the same as 82%. So you need to find the log to the base of a half of 0.82. (And then take a third of that). Follow?

OpenStudy (anonymous):

(And then take the reciprocal of that.)

OpenStudy (dumbcow):

\[P _{t} = P _{0}e ^{-kt}\] First find k using P and t given assume P is percent of drug in bloodstream P0 is initial condition, 100% in bloodstream \[0.82 = e ^{-3k}\] take ln both sides \[k = \frac{\ln .82}{-3}\] Now we needto solve for t when 50% of drug is in bloodstream(half-life) \[0.5 = e ^{\frac{\ln .82}{3}t}\] take ln of both sides \[t = \ln .5 * \frac{3}{\ln .82} = 10.48\]